# Pushing the Accuracy Limit of Foundation Neural Network Models with Quantum Monte Carlo Forces and Path Integrals

### v1

@@
Pushing the Accuracy Limit of Foundation Neural Network Models with Quantum Monte Carlo Forces and Path Integrals
Anouar Benali Qubit Pharmaceuticals Inc,
Boston, MA 02116, USA
   Thomas Plé Laboratoire de Chimie
Théorique, Sorbonne Université,
UMR 7616 CNRS,
75005 Paris, France
   Olivier Adjoua Laboratoire de Chimie
Théorique, Sorbonne Université,
UMR 7616 CNRS,
75005 Paris, France
   Valay Agarawal Department of Chemistry,
University of Chicago,
Chicago, IL 60637, USA
   Thomas Applencourt Argonne Leadership Computing Facility,
Argonne National Laboratory,
Argonne, IL 60439, USA
   Marharyta Blazhynska Laboratoire de Chimie Théorique,
Sorbonne Université,
UMR 7616 CNRS,
75005 Paris, France
   Raymond Clay III HEDP Theory Department,
Sandia National Laboratories,
Albuquerque, NM 87185, USA
   Kevin Gasperich Qubit Pharmaceuticals Inc,
Boston, MA 02116, USA
   Khalid Hossain Argonne Leadership Computing Facility,
Argonne National Laboratory,
Argonne, IL 60439, USA
   Jeongnim Kim Intel Corporation,
Hillsboro, OR 987124, USA
   Christopher Knight Computational Science Division,
Argonne National Laboratory,
Argonne, IL 60439, USA
   Jaron T. Krogel Materials Science and Technology Division,
Oak Ridge National Laboratory,
Oak Ridge, TN 37831, USA
   Yvon Maday Sorbonne Université,
Université Paris Cité,
CNRS, INRIA, LJLL UMR 7598,
Paris 75005, France
   Maxime Maria Université de Limoges,
XLIM UMR CNRS,
87060 Limoges, France
   Mathieu Montes Conservatoire National des Arts et Métiers,
GBCM, 75003 Paris, France
   Ye Luo Computational Science Division,
Argonne National Laboratory,
Lemont, IL 60439, USA
   Evgeny Posenitskiy Qubit Pharmaceuticals,
75014 Paris, France
   Corentin Villot Qubit Pharmaceuticals,
75014 Paris, France
   Venkat Vishwanath Argonne Leadeship Computing
Facility, Argonne National Laboratory,
Argonne, IL 60439, USA
   Louis Lagardère Laboratoire de Chimie Théorique,
Sorbonne Université,
UMR 7616 CNRS
and Qubit Pharmaceuticals, Paris, France
   Jean-Philip Piquemal Laboratoire de Chimie Théorique,
Sorbonne Université,
UMR 7616 CNRS
and Qubit Pharmaceuticals, Paris, France
Email: jean-philip.piquemal@sorbonne-universite.fr
@@ Abstract
We propose an end-to-end integrated strategy for the production of highly accurate quantum chemistry (QC) synthetic datasets aimed at deriving atomistic Foundation Machine Learning (ML) Models. We first present a GPU-accelerated QC database generation Exascale protocol able to produce the required energies and forces. A ”Jacob’s Ladder” approach leverages computationally-optimized layers of massively parallel high performance software with increasing accuracy to compute: i) Density Functional Theory (DFT); ii) Quantum Monte Carlo (QMC); iii) Selected Configuration Interaction (sCI), within large volumes and optimized time-to-solution performances. Handling this ambitious computational pipeline would be impossible without exascale computing resources, particularly for the notoriously difficult and computationally intensive calculation of QMC forces and for the combination of multi-determinant QMC energies and forces using selected CI wavefunctions methodologies. To our knowledge, this is the first time that such quantities are computed at such scale. We combine these data with the FeNNix-Bio-1 foundation ML model to bridge the gap between highly accurate QC calculations and condensed-phase Molecular Dynamics (MD). We demonstrate stable multi-ns simulations using the resulting beyond DFT accuracy fully reactive model coupled to full path integrals adaptive sampling quantum dynamics. A complete 1 million-atom plant virus solvated structure, including its full genetic material, is simulated using Ring-Polymer MD quantum dynamics along as its response to acidification under physiological NaCl concentrations. These new capabilities open the door to the possibility to monitor bond breaking/creation and proton transfers chemical interactions taking place in biosystems allowing us to reach a deeper understanding of their complex internal machinery.
1 Justification for Prize
  • Record Quantum Monte Carlo and Selected Configuration Interaction (energies and forces) computations in terms of number of calculations: 20,251 QMC (energies+forces), 20,338 QMC (energies-only), 2000 s-CI (energies+forces); and achieved accuracy. Double precision performance of 124/62/52 PFLOPS on 2048 nodes of Aurora (12,288 GPUs), representing approximately 60%/30% and 25% of theoretical peak performance for energy calculations, force calculations, and multideterminant force calculations, respectively.
  • For a 32-million-particule (1-million atom X 32 beads) Satellite Tobacco Virus system (STMV) Ring Polymer quantum molecular dynamics simulation with a foundation neural network model (using project quantum chemistry data) while using multi-replica via an unsupervised adaptive sampling strategy, time-to-solution of 0.44ns/day (128 nodes, 1024 GPUs, mixte precision) on Jean Zay. Establish new state-of-the-art for foundation model quantum dynamics simulation of biological system. Simulations captured the STMV response to acidification under physiological NaCl concentrations for the first time.
2 Performance Attributes
@@
Type of method used (2) Neural Network potential Molecular Dynamics
@@
System scale Measured on full systems / extrapolated times
@@
3 Overview of the problem
Simulating the structural dynamics of biological systems at full quantum accuracy is a daunting task requiring precise modeling of the Born-Oppenheimer electronic potential energy surface while including nuclear quantum effects in the dynamics to accurately reproduce condensed phase properties. Up to now, achieving such high-fidelity simulations has been prevented by the computational cost of first-principles quantum approaches and remains a Graal for the molecular simulation community. In this context, neural networks potentials have started to gain speed and precision offering a clear alternative to ab initio molecular dynamics (AIMD) and recently even to force fields simulations thanks to high performance computing. Thus, the field evolves very fast, and larger foundation models dedicated to atomistic Molecular Simulation have been recently introduced. These single general-purpose ML potentials have various capabilities and can be applied to various types of applications, efficiently replacing quantum mechanical approaches. However, in practice, such ML approaches cannot perform computations at an accuracy better than the models’ reference data. Therefore, foundation models rely heavily on the quality and the quantity of their reference synthetic data. The choice of the quantum chemistry approaches and their capability to approximate the Schrodinger equation therefore have a considerable impact on the performance of the model. Additionally, on the molecular dynamics side, since foundation models can be seen as Born-Oppenheimer potential surfaces generators, it is critical to go beyond classical dynamics. Indeed, it is paramount to be able to include the effects of nuclei to perform quantum dynamics to access real condensed-phase properties. Overall, the rise of exascale computing offers the possibility of designing a new, multi-level, pipeline for the generation of accurate quantum chemistry data that would strongly impact future fondation machine learning model developments.
4 Current State of the art
4.0.1 Quantum Chemistry Datasets
@@
By contrast, Quantum Monte Carlo (QMC) methods[73], and particularly Diffusion Monte Carlo (DMC)[74, 27], have proven capable of near-exact ground-state energies across wide chemical spaces[28, 85, 84, 48, 83, 61, 11, 10, 47, 23]. Large-scale QMC studies focused on benchmark energies[12, 44], but until recently lacked an efficient pathway to obtain forces, rendering them less suitable for force field generation. Computing forces in Diffusion Monte Carlo (DMC) remains inherently challenging due to the fixed-node approximation and the intrinsic stochastic nature of the algorithm. Fortunately, over the past several years, a variety of technical[26], mathematical[96], and performance improvements[25] to the original Zero-Variance Zero-Bias force estimator framework of Assaraf and Caffarel[4] have made QMC forces usable and trustworthy for large scale applications.
@@
4.1 Foundation Machine Learning Models for atomistic molecular dynamics Simulations
The application of Neural Networks to molecular dynamics is recent and started with the availability of suitable neural architecture  [9, 6]. Since then, numerous neural network potentials (NNPs) have been proposed [20, 14, 9, 81, 88, 78, 77, 31, 55, 104, 101, 71, 94, 30, 8, 62, 34, 35, 21, 105, 66, 46, 100] while their high performance implementations reached a mature stage [86, 41, 39] and as specialized libraries are now available to NNPs developers  [79, 22, 102, 67]. Very recently, in link with the breakthroughs of the AlphaFold [42] and RoseTTAFold [5] protein structure foundation models, a major shift in the developments started to reshape molecular simulation with introduction of the MACE-MP-0 foundation model. [7]. As explained by its authors, MACE is a single general-purpose ML potential, designed for material science atomistic simulations. Its capabilities range from single molecule to full materials simulations, and enables modeling of chemical reactivity. Beyond solids, its extended transferability allowed it to be tested even on small proteins where stable molecular dynamics could be obtained. However, since MACE and its associated neural architecture are primarily designed for materials, its applicability to condensed phase simulations of large complex biological molecules remains limited both in term of accuracy and computational performances. For these reasons, some of us recently introduced the FeNNix-Bio1 Foundation Model[68]. Note that its associated paper can be seen as the machine learning counterpart of the present Exascale computing report as both projects were launched together and as the FeNNix-Bio1 development used preliminary data extracted from the present quantum chemistry pipeline. Thus, it is able to perform atomistic condensed phase molecular dynamics simulations of biological systems while allowing for reactivity modeling. FeNNix-Bio1 is fast and accurate, being able to address simultaneously the issues of the computational requirements generated by the biological timescales and the difficult transferability of charged-species, that are numerous in Chemistry and Biology. Finally, FeNNix-Bio1 is also able to compute meaningful condensed-phase properties since it explicitly takes into account nuclear quantum effects in its (quantum) dynamics. [18, 60, 69]. This allows the foundation model to predict accurate condensed phase properties such as free energies, solvent properties, etc. .
5 Innovations
5.1 Dataset Generation: High Performance Quantum Chemistry Pipeline
@@
5.1.1 The SPICE2(+)-ccECP dataset
@@
5.1.2 Density Functional Theory
@@
5.2 Quantum Monte Carlo
QMC techniques are stochastic methods that solve the many-body Schrödinger equation with high accuracy using a limited but controlled number of approximations. By explicitly including many-body electronic interactions, these methods achieve chemical accuracy while maintaining favorable scaling compared to traditional high-level quantum chemistry approaches. In this work, we leverage the extensive refactoring of QMCPACK that has been specifically optimized for exascale computing platforms over the past two decades. Building on previous optimizations that targeted material science applications, we extended and enhanced the Linear Combination of Atomic Orbitals (LCAO) branch of QMCPACK with new optimizations and data layouts specifically designed for GPU acceleration in molecular applications. These enhancements enable unprecedented scaling on Aurora’s architecture for complex molecular systems. Among the various QMC approaches, we employed Diffusion Monte Carlo (DMC) due to its ability to recover a significant portion of the correlation energy while maintaining favorable scaling with system size. DMC solves the Schrödinger equation in imaginary time $\tau=it$, ensuring that any initial state $|\psi\rangle$ not orthogonal to the ground state $|\phi_{0}\rangle$ will converge to the ground state in the long-time limit:
@@ 5.2 Quantum Monte Carlo
5.2.1 QMC Energy and Wavefunction Evaluation
@@
5.2.2 Single Determinant QMC Forces
@@ 5.2.3 Multideterminant QMC Forces
5.2.4 Selected Configuration Interaction
@@
Our sCI code is only several months old and is still undergoing rapid development. There are many more algorithmic improvements underway and planned for the near future, but it is already capable of leveraging GPUs, and for this work it was able to generate 1.4M-determinant wavefunctions for 2000 molecules using approximately 1k node-hours, which represents less than 1% of the total computational cost of the full {DFT→sCI→VMC/DMC} workflow per molecule. So even at a very early stage of development, with many planned improvements that will increase the capability of the code, we are still able to greatly improve our nodal surface quality (by adding multideterminant expansions) without adding any significant cost in performing the sCI calculations.
@@
5.3 Performance Analysis and Benchmarks
@@
5.3.1 Single-Determinant QMC Performance
@@
This stark performance gap highlighted the need for fundamental algorithm redesign, specifically addressing GPU kernel overhead, parallel efficiency, and memory optimization as discussed in Sections 5.2.1 and 5.2.2.
@@
5.3.2 Resource Allocation Strategy
@@ 5.3.3 Multideterminant QMC Performance
5.4 Large Scale Quantum Dynamics Simulations with a ”beyond DFT” Foundation Neural Network Model
@@
Using the QMC and sCI data produced in the project, we improved an early version of the FeNNix-Bio1 model [68] using transfer machine learning via the FeNNol library[67] to obtain a ”beyond DFT” improved approach. We then coupled the improved model to the GPU-accelerated massively parallel Tinker-HP molecular dynamics package [49, 1] and leverage simultaneously its Deep-HP (neural networks), Quantum-HP (quantum nuclear effects) modules while using enhanced sampling via non-supervised adaptive sampling molecular dynamics. We tested the new approach on both Jean-Zay (H100) and Aurora(Xe) machines on the full structure of the Satellite Tobacco Virus, a 1-million atom virus. The simulation encompassed 32-million particles (1-million atom X 32 beads) since it leverages Ring Polymer quantum molecular dynamics simulation. We extended the conformational space sampling with adaptive sampling replica to decompose the global exploration problem into a set of separate MD trajectories. These latter can be restarted thanks to a selective process in order to achieve sufficient phase-space sampling.
@@
    We coupled our newly developed JAX-based [16] FeNNoL library [67] that allows for the construction and inference of atomistic machine learning potentials such as FeNNix-Bio1[68] to the Tinker-HP molecular dynamics package[49, 1] via the Deep-HP interface[39]. By doing so, we leverage the Ring Polymer Molecular Dynamics (RPMD)[36] capabilities of Tinker-HP (Quantum-HP module [70]) in combination with FeNNix-Bio1 in order to recover quantum dynamics and statistics of the Satellite Tobacco Virus (STMV) that we previously studied [15] using polarizable molecular dynamics. Accounting for these has been shown to be mandatory to recover correct properties in that context. Such simulations leverage two levels of parallelism: one at the beads level and a second one within each bead. The choice of the STMV was motivated as it is a model system for studying viral assembly. Moreover, since it is includes a single-stranded RNA genome located within an icosahedral capsid, it is also a good model to study the structural dynamics stability of viruses[15]. Finally, to push further the sampling of the conformational space, we added on top of the discussed framework an additional layer of parallelism using automatized (unsupervised) adaptive sampling following described previously[40]. Here multiple iterations of independent molecular dynamics simulations replica are launched. The framework is an iterative process and the initial conditions at each iteration are based on the results of previous iteration steps. The selection of the initial structures at each iteration follows an adaptive, fully automated, procedure enhancing the exploration of a low-dimensional space of slow variables. This unsupervised selection step has the advantage of suppressing the problem of the choice of the initial collective variable at the beginning of the simulation. In practice the procedure maximize the conformational space exploration by penalizing areas that have already been extensively visited. The choice of the number of replica is only limited by the number of available GPUs, making adaptive sampling particularly suited for exascale computing (see section of reference [40] for a detailed discussion. On Jean-Zay, each of the bead was parallelized on 8 H100 GPUs for a total of 8x32=256 H100s. Exploiting parallelism within each bead is mandatory to fit in the memory of each GPU thanks to memory distribution. All in all, we obtain 0.11 ns/day (mixte precision) per single adaptive sampling replica, making production of tens of nanoseconds per day of such quantum dynamics simulation accessible through adaptive sampling given enough resources. To our knowledge, achieving such level of quantum dynamics is unprecedented. The Jean Zay machine having 1280 GPUs, we were able to launch 4 simultaneous replica leveraging 1024 GPUs, producing therefore 0.44 ns/day of adaptive quantum dynamics. This motivated the portage to Exascale on Aurora to increase time to solution thanks to the large availability of computing nodes.

### v2

@@ Abstract
We propose an end-to-end integrated strategy to produce highly accurate quantum chemistry (QC) synthetic datasets (energies and forces) aimed at deriving Foundation Machine Learning models for molecular simulation. Starting from Density Functional Theory (DFT), a ”Jacob’s Ladder” approach leverages computationally-optimized layers of massively parallel GPU-accelerated software with increasing accuracy. Thanks to Exascale, this is the first time that the computationally intensive calculation of Diffusion Quantum Monte Carlo (QMC) forces, and the combination of multi-determinant QMC energies and forces with selected-CI wavefunctions, are computed at such scale at the complete basis-set-limit. To bridge the gap between accurate QC and condensed-phase molecular dynamics, we leverage transfer learning to improve the FeNNix-Bio1 DFT-based foundation model. The resulting approach is coupled to path integrals adaptive sampling quantum dynamics to perform nanosecond reactive simulations at unprecedented accuracy. These results demonstrate the promise of Exascale to deepen our understanding of the inner machinery of complex biosystems.
I Justification for Prize
  • Record Quantum Monte Carlo (Diffusion Monte Carlo - DMC) and Multideterminant DMC (from selected CI nodes (sCI)) computations were performed, including 20,338 QMC (energies-only), 20,251 QMC (energies + forces), and 2,000 multideterminant Diffusion Monte Carlo (energies + forces) calculations using determinants from sCI. These achieved double-precision performance of 124, 62, and 52 PFLOPS, respectively, on 2,048 nodes of Aurora (12,288 GPUs), corresponding to approximately 60%, 30%, and 25% of the system’s theoretical peak performance for energies-only, forces, and multideterminant forces evaluations. All QMC simulations were performed in real space and are thus free from basis-set errors.
  • For a 32-million-particule (1-million atom X 32 beads) Satellite Tobacco Virus system (STMV) Ring Polymer quantum molecular dynamics simulation with a foundation machine learning (ML) model (using the present project high accuracy quantum chemistry data) while using multi-replica via an unsupervised adaptive sampling strategy, time-to-solution of 0.44ns/day (4 replica, 128 nodes, 1024 GPUs, mixed precision on Jean Zay or 15 replica, 480 nodes, 5760 GPUs, mixed precision on Aurora). Establish new state-of-the-art for reactive foundation model atomistic quantum dynamics simulation of biological systems.
II Performance Attributes
@@
Type of method used (2) ML potential molecular dynamics
@@
System scale Measured on full system
@@
III Overview of the problem
Simulating the structural dynamics of biological systems at full quantum accuracy is a daunting task requiring precise modeling of the Born-Oppenheimer electronic potential energy surface while including nuclear quantum effects in the dynamics to accurately reproduce condensed phase properties. Up to now, achieving such high-fidelity simulations has been prevented by the computational cost of first-principles quantum approaches that can only reach the required accuracy on very small systems and remains therefore a Graal for the molecular simulation community. In this context, neural networks potentials have started to gain speed and precision offering a clear alternative to ab initio molecular dynamics (AIMD) and recently even to force fields simulations thanks to high performance computing. Thus, the field evolves very fast, and larger foundation models dedicated to atomistic Molecular Simulation have been recently introduced. These single general-purpose ML potentials have various capabilities and can be applied to various types of applications, efficiently replacing quantum mechanical approaches. However, in practice, such ML approaches cannot perform computations at an accuracy better than the models’ reference data. Therefore, foundation models rely heavily on the quality and the quantity of their reference synthetic data. The choice of the quantum chemistry approaches and their capability to approximate the Schrodinger equation therefore have a considerable impact on the performance of the model. However, in principle, with enough extremely accurate small molecules quantum chemistry data, one could transfer such accuracy to very large systems. Additionally, on the molecular dynamics side, since foundation models can be seen as Born-Oppenheimer potential surfaces generators, it is critical to go beyond classical dynamics. Indeed, in this precise context where no ”condensed-phase parametrization” occurs like in usual force fields, it becomes paramount to include the effects of nuclei via quantum dynamics in order to predict accurate condensed-phase properties. Overall, the rise of exascale computing offers the possibility of designing a new, multi-level, pipeline for the generation of accurate quantum chemistry data that would strongly impact future foundation machine learning model developments. Indeed, ultimately, these latter could become capable of performing large scale, chemically accurate, molecular simulations.
IV Current State of the art
IV-1 Quantum Chemistry Datasets
@@
By contrast, Quantum Monte Carlo (QMC) methods[75], and particularly Diffusion Monte Carlo (DMC)[76, 27], have proven capable of near-exact ground-state energies across wide chemical spaces[28, 87, 86, 49, 85, 63, 11, 10, 47, 23]. Large-scale QMC studies focused on benchmark energies[12, 44], but until recently lacked an efficient pathway to obtain forces, rendering them less suitable for force field generation. Computing forces in Diffusion Monte Carlo (DMC) remains inherently challenging due to the fixed-node approximation and the intrinsic stochastic nature of the algorithm. Fortunately, over the past several years, a variety of technical[26], mathematical[99], and performance improvements[25] to the original Zero-Variance Zero-Bias force estimator framework of Assaraf and Caffarel[4] have made QMC forces usable and trustworthy for large-scale applications.
Unlike traditional quantum chemistry methods, which rely on finite basis sets and must extrapolate to the complete basis set (CBS) limit or depend on error cancellation and ad hoc corrections, QMC—particularly DMC—operates directly in real space. As a result, it is inherently free from basis set incompleteness error and provides results effectively in the infinite-basis limit without the need for basis-set tricks or extrapolation schemes.
@@
IV-A Foundation Machine Learning Models for atomistic molecular dynamics Simulations
The application of Neural Networks to molecular dynamics is recent and started with the availability of suitable neural architecture  [9, 6]. Since then, numerous neural network potentials (NNPs) have been proposed [20, 14, 9, 83, 90, 80, 79, 31, 56, 108, 104, 73, 97, 30, 8, 64, 34, 35, 21, 109, 68, 46, 103] while their high performance implementations reached a mature stage [88, 41, 39] and as specialized libraries are now available to NNPs developers  [81, 22, 105, 69]. Very recently, in link with the breakthroughs of the AlphaFold [42] and RoseTTAFold [5] protein structure foundation models, a major shift in the developments started to reshape molecular simulation with the introduction of the MACE-MP-0 foundation model for molecular dynamics. [7]. As explained by its authors, MACE is a single general-purpose ML potential, designed for material science atomistic simulations. Its capabilities range from single molecule to full materials simulations, and enables modeling of chemical reactivity. Beyond solids, its extended transferability allowed it to be tested even on small proteins where stable molecular dynamics could be obtained. However, since MACE and its associated neural architecture are primarily designed for materials, its applicability to condensed phase simulations of large complex biological molecules remains limited both in term of accuracy and computational performances. For these reasons, some of us recently introduced the FeNNix-Bio1 Foundation Model[72] for atomistic molecular dynamics simulations. Note that its associated paper can be seen as the machine learning counterpart of the present Exascale computing report as both projects were launched together and the FeNNix-Bio1 development used some preliminary data extracted from the present quantum chemistry pipeline. Thus, it is able to perform atomistic condensed phase molecular dynamics simulations of biological systems while allowing for reactivity modeling. FeNNix-Bio1 is fast and accurate, being able to address simultaneously the issues of the computational requirements generated by the biological timescales and the difficult transferability of charged-species, that are numerous in Chemistry and Biology. Finally, FeNNix-Bio1 is also able to compute meaningful condensed-phase properties since it explicitly takes into account nuclear quantum effects in quantum dynamics. [18, 62, 70]. This allows the foundation model to predict accurate condensed phase properties such as free energies, solvent properties, etc. .
V Innovations
V-A Dataset Generation: High Performance Quantum Chemistry Pipeline
@@
V-A1 The SPICE2(+)-ccECP dataset
@@
V-A2 Density Functional Theory
@@
V-B Quantum Monte Carlo
QMC techniques are stochastic methods that solve the many-body Schrödinger equation with high accuracy using a limited but controlled number of approximations. By explicitly including many-body electronic interactions, these methods achieve chemical accuracy while maintaining favorable scaling compared to traditional high-level quantum chemistry approaches. In this work, we leverage the extensive refactoring of QMCPACK that has been specifically optimized for exascale computing platforms over the past two decades. Building on previous optimizations that targeted material science applications, we extended and enhanced the Linear Combination of Atomic Orbitals (LCAO) branch of QMCPACK with new optimizations and data layouts specifically designed for GPU acceleration in molecular applications. These enhancements enable unprecedented scaling on Aurora’s architecture for complex molecular systems. Among the various QMC approaches, we employed Diffusion Monte Carlo (DMC) due to its ability to recover a significant portion of the correlation energy while maintaining favorable scaling with system size. Importantly, DMC operates directly in real space, meaning it does not rely on a finite basis set. As a result, DMC is free from basis set incompleteness error and can be viewed as operating in the infinite-basis limit by construction. DMC solves the Schrödinger equation in imaginary time $\tau=it$, ensuring that any initial state $|\psi\rangle$ not orthogonal to the ground state $|\phi_{0}\rangle$ will converge to the ground state in the long-time limit:
@@ V-B Quantum Monte Carlo
V-B1 QMC Energy and Wavefunction Evaluation
@@
V-B2 Single Determinant QMC Forces
@@ V-B3 Multideterminant QMC Forces
V-B4 Selected Configuration Interaction
@@
Our sCI code is only several months old and is still undergoing rapid development. There are many more algorithmic improvements underway and planned for the near future, but it is already capable of leveraging GPUs, and for this work it was able to generate 1.4M-determinant wavefunctions for 2000 molecules using approximately 1k node-hours, which represents less than 1% of the total computational cost of the full {DFT→sCI→VMC/DMC} workflow per molecule. So even at a very early stage of development, with many planned improvements that will increase the capability of the code, we are still able to greatly improve our nodal surface quality (by adding multideterminant expansions) without adding any significant cost in performing the sCI calculations. Crucially, because we ultimately project these wavefunctions with DMC, we eliminate the limitations of the finite basis used in sCI. The final energies and forces are obtained in real space and are thus free from basis set incompleteness error, making the approach both efficient and systematically improvable.
@@
V-B5 QMC Submission Engine on Aurora
QMCPACK calculations on Aurora were facilitated by a custom two-level job submission engine built on top of the Nexus[48] workflow system. To enable to maximize computational efficiency, molecular systems were batched together by electron count similarity as this supports greater execution uniformity among batched molecules. This top level parallelism, across batches, was complemented by further separation of molecules into sub-ensembles, each of which ran under a single top-level QMCPACK MPI process to maximize robustness of the simulations, providing a simple form of fault tolerance. The balance between these levels, as well as assignment of individual molecules to selected node resources, was optimized to simulaneously maximize job submission size, fault tolerance, and computational efficiency.
V-C Performance Analysis and Benchmarks
@@
V-C1 Single-Determinant QMC Performance
@@
This stark performance gap highlighted the need for fundamental algorithm redesign, specifically addressing GPU kernel overhead, parallel efficiency, and memory optimization as discussed in Sections V-B1 and V-B2.
@@
V-C2 Resource Allocation Strategy
@@ V-C3 Multideterminant QMC Performance
V-D Large Scale Quantum Dynamics Simulations with a ”beyond DFT” Foundation Neural Network Model
@@
Using the complete basis-set QMC data produced in the project, we improved an early version of the FeNNix-Bio1 model [72] using transfer machine learning via the FeNNol library[69] to obtain a ”beyond DFT” improved approach. We then coupled the improved model to the GPU-accelerated massively parallel Tinker-HP molecular dynamics package [50, 1] and leverage simultaneously its Deep-HP (neural networks), Quantum-HP (quantum nuclear effects) modules while using enhanced sampling via non-supervised adaptive sampling molecular dynamics. We tested the new approach on both Jean-Zay (H100) and Aurora(Xe) machines on the full structure of the Satellite Tobacco Virus, a 1-million atom virus. The simulation encompassed 32-million particles (1-million atom X 32 beads) since it leverages Ring Polymer quantum molecular dynamics simulation. We extended the conformational space sampling with adaptive sampling replica to decompose the global exploration problem into a set of separate MD trajectories. These latter can be restarted thanks to a selective process in order to achieve sufficient phase-space sampling.
@@
  • Leveraging Transfer Learning to Incorporate High Accuracy Quantum Chemistry Data into FeNNix-Bio1
    As a proof of concept, in order to improve further the quality of the DFT-based FeNNix-Bio1 model towards a ”beyond DFT” model, we leveraged transfer learning,[92] to include the high-quality, complete basis-set, DMC energies and forces present in the SPICE2(+)-ccECP dataset. In practice, starting from the model trained on the DFT part of the dataset, we reoptimize the energy heads (freezing the parameters of the embedding), on the QMC energies and forces with the same loss function as in the original training. As is common practice, we started with a smaller learning rate of $10^{-5}$ using the adabelief optimizer[107].
@@
    We coupled our newly developed JAX-based [16] FeNNoL library [69] that allows for the construction and inference of atomistic machine learning potentials such as FeNNix-Bio1[72] to the Tinker-HP molecular dynamics package[50, 1] via the Deep-HP interface[39]. By doing so, we leverage the Ring Polymer Molecular Dynamics (RPMD)[36] capabilities of Tinker-HP (Quantum-HP module [71]) in combination with FeNNix-Bio1 in order to recover quantum dynamics and statistics of the Satellite Tobacco Virus (STMV) that we previously studied [15] using polarizable molecular dynamics. Indeed RPMD approximates the Kubo-transformed time correlation function and can be used efficiently for determining accurate dynamical properties of condensed-phase molecular systems. It allows to use classical statistical physics to mimic quantum dynamics via fictitious ring polymer which consist in n copies (or beads) of the system particles connected by harmonic springs. The interested reader can refer to ref.[71] for details about RPMD and its implementation. Accounting for these has been shown to be mandatory to recover correct properties in that context and up to now FeNNix-BIo1 only resorted to more approximated methods such as adQTB[62] to mimic quantum dynamics. If adQTB is extremely fast, i.e. roughly similar computational cost as classical dynamics, and well suited for most of the FeNNix-BIo1 applications [72], the present RPMD implementation allows to reach an unprecedented accuracy on the FeNNiX-Bio1 quantum dynamics leading to an improved control on the chemical reactions, notably proton transfers. To ensure full accuracy at room temperature, we used 32 beads per atom in RPMD simulations. These latter leverage two levels of parallelism: one at the beads level and a second one within each bead. The choice of the STMV was motivated as it is a model system for studying viral assembly. Moreover, since it is includes a single-stranded RNA genome located within an icosahedral capsid, it is also a good model to study the structural dynamics stability of viruses[15]. Finally, to push further the sampling of the conformational space, we added on top of the discussed framework an additional layer of parallelism using automatized (unsupervised) adaptive sampling following described previously[40]. Here multiple iterations of independent molecular dynamics simulations replica are launched to ”mine” the conformational space. The framework is an iterative process and the initial conditions at each iteration are based on the results of previous iteration steps. The selection of the initial structures at each iteration follows an adaptive, fully automated, procedure enhancing the exploration of a low-dimensional space of slow variables. This unsupervised selection step has the advantage of suppressing the problem of the choice of the initial collective variable at the beginning of the simulation. In practice the procedure maximize the conformational space exploration by penalizing areas that have already been extensively visited. The choice of the number of replica is only limited by the number of available GPUs, making adaptive sampling particularly suited for exascale computing (see section of reference [40] for a detailed discussion. On Jean-Zay, each of the bead was parallelized on 8 H100 GPUs for a total of 8x32=256 H100s. Exploiting parallelism within each bead is mandatory to fit in the memory of each GPU thanks to memory distribution. All in all, we obtain 0.11 ns/day (mixed precision) per single adaptive sampling replica, making production of tens of nanoseconds per day of such quantum dynamics simulation accessible through adaptive sampling given enough resources. To our knowledge, achieving such level of quantum dynamics with a foundation model and grounded on extremely accurate quantum chemistry dataset on such a large biosystem is unprecedented. The Jean Zay machine having 1280 GPUs, we were able to launch 4 simultaneous replica leveraging 1024 GPUs, producing therefore 0.44 ns/day of adaptive quantum dynamics. Note that such adaptive setup can also be used sequentially to extend the number of replica that can be launched one after another enabling to accumulate simulation time. Overall, this computing scheme is therefore particularly suited for large exascale machine and further H100 optimization of the combined computing layers should possibly enable to reach the 1 ns/day threshold.